Find the second derivative of 3y^4+x^7=5x

1
Expert's answer
2021-12-31T08:55:03-0500

3y⁴+x⁷=5x

Differentiating with respect to x

12y³"\frac{dy}{dx}+7x⁶ = 5"

=> 12y³"\frac{dy}{dx}" = 5 - 7x⁶

=> "\frac{dy}{dx}" = "\frac{5-7x⁶}{12y³}" ••••••••(1)

Now 12y³"\frac{dy}{dx}+7x⁶ = 5"

Differentiating again with respect to x

"12y³\frac{d²y}{dx²}+ 36y²{(\frac{dy}{dx})}^2" + 42x⁵ = 0

=> "2y³\frac{d²y}{dx²}+ 6y²{(\frac{dy}{dx})}^2" + 7x⁵ = 0

=>

"2y³\frac{d²y}{dx²}+ 6y²{(\frac{5-7x⁶}{12y³})}^2+7x⁵=0"

=> "2y³\frac{d²y}{dx²}= -6y²{(\frac{5-7x⁶}{12y³})}^2-7x⁵"

=> "\frac{d²y}{dx²}= \frac{ -6y²{(\frac{5-7x⁶}{12y³})}^2-7x⁵}{2y³}"

=> "\frac{d²y}{dx²}=- \frac{ 6y²{(\frac{5-7x⁶}{12y³})}^2+7x⁵}{2y³}"

=> "\frac{d²y}{dx²}=- \frac{ 6y²{(\frac{(5-7x⁶)²}{144y⁶})}+7x⁵}{2y³}"

=> "\frac{d²y}{dx²}=- \frac{ {(\frac{(5-7x⁶)²}{24y⁴})}+7x⁵}{2y³}"

=> "\frac{d²y}{dx²}=- \frac{ (5-7x⁶)²+168x⁵y⁴}{48y⁷}"